OpenStudy (anonymous):

What is the limit of (2-x)/(x^2-4) as x approaches 2?

Directrix (directrix):

-1/4 See if you agree with the work in the attached file.

OpenStudy (anonymous):

Is one allowed to factor out a (-1) like that?

Directrix (directrix):

Yes. The -1 was factored out of the numerator only. I learned this as the "three signs of a fraction." Example: -(1/2) = (-1)/2 = 1/(-2) Note: -(1/2) is never equal to (-1)/(-2) {which would be positive 1/2}

OpenStudy (anonymous):

Thank you very much!